# 360VOTS: Visual Object Tracking and Segmentation in Omnidirectional Videos

### v1

@@ I Introduction
Furthermore, we present 360VOTS, a challenging benchmark dataset to facilitate research and evaluation on omnidirectional tracking and segmentation, including 360VOT [23] and 360VOS datasets. 360VOT provides dense (rotated) BBox and BFoV annotations of 120 test sequences. Inheriting parts of sequences from the 360VOT dataset, 360VOS incorporates additional 191 sequences and 30 tracking object categories. This results in a large-scale dataset consisting of a total of 290 sequences in 62 categories. Importantly, 360VOS provides dense pixel-wise annotations as ground truth and assigns 170 sequences as the training set. These pixel-wise annotations are utilized to obtain unbiased (rotated) BBoxes and BFoVs ground truth, ensuring comprehensive evaluation and comparison of different approaches. To accurately evaluate the performance of state-of-the-art (SOTA) approaches on omnidirectional tracking and segmentation, we develop new metrics tailored for 360°$\degree$ images. These metrics encompass dual success, dual (normalized) precision, and angle precision to assess tracking performance, as well as spherical region similarity and spherical contour accuracy to evaluate segmentation performance. Moreover, we exploit the 360VOS training set to retrain the tracker and apply our 360 tracking framework to achieve new SOTA results. With the introduction of new representations, datasets, metrics, and baselines, we believe 360VOTS serves as a vital resource to foster development in omnidirectional visual object tracking and segmentation.
@@
  • We benchmark 20 tracking algorithms and 16 segmentation algorithms on 360VOT and 360VOS respectively, and develop new baselines for future comparisons.
@@ II-A Benchmarks for Visual Object Tracking
ALOV300 [24] is a sparse benchmark introducing 152K frames and 16K annotations, while UAVDT [29] focuses on UAV scenarios and has 100 videos. TrackingNet [31] is a large-scale dataset collecting more than 14M frames based on the YT-BB dataset [34]. As YT-BB only provides fine-grained annotations at 1 FPS, they explored a tracker to densify the annotations without further manual refinement. OxUvA [30] targets long-term tracking by constructing 337 video sequences, but each video only has 30 frames annotated.
One of the first dense BBox benchmarks is OTB100 [11] which is extended from OTB50 [35] and has 100 sequences. NUS-PRO [25] takes the feature of occlusion-level annotation and provides 365 sequences, while TC128 [26] researches the chromatic information in visual tracking. UAV123 [12] and DTB70 [27] offer 123 and 70 aerial videos of rigid objects and humans in various scenes. NfS [28] consists of more than 380K frames captured at 240 FPS studying higher frame rate tracking, while LaSOT [14] is a large-scale and category balanced benchmark of premium quality. GOT-10k [15] provides about 1.5M annotations and 84 classes of objects, aiming at generic object tracking. The annual tracking challenge VOT [13] offered 62 sequences and 20K frames in 2022. A more recent benchmark TOTB [32] mainly focuses on transparent object tracking. TREK-150 [33] introduces 150 sequences of tracking in First Person Vision (FPV) with the interaction between the person and the target object.
@@ II-B Benchmarks for Video Object Segmentation
There are several benchmarks for video object segmentation proposed in the last decade as described in Table I. FBMS-59[16] acknowledges the importance of motion information and concentrates on motion segmentation, providing a total of 13,860 frames across 59 sequences. DAVIS[18] is the first dataset designed for video object segmentation (VOS), annotating 376 classes of objects on each frame of 120 short videos. Leveraging the diverse scenarios and photography conditions on YouTube videos, YouTube-VOS[19] introduces a large-scale VOS dataset with more than 540K frames and contains 3471 and 508 sequences as the training set and the test set separately. Each video clip has a duration of 3 to 6 seconds and is annotated every 5 frames. One of the most recent datasets is LVOS[17] which is the first long-term VOS benchmark dataset with an average length of 560 densely labeled frames. In contrast, our 360VOS provides pixel-wise masks of omnidirectional images for both training and testing.
II-C Adapting Object Tracking Schemes in Segmentation
To guarantee high tracking speed, the trackers for single object tracking generally crop the image and search for the target in small local regions. The tracking scheme is vital in selecting searching regions and interpreting network predictions over sequences in the inference phase. A compatible tracking inference scheme can enhance tracking performance. For example, DaSiamRPN [36] explored a local-to-global searching strategy for long-term tracking. SiamX [37] proposed an adaptive inference scheme to prevent tracking loss and realized fast target re-localization. Similarly, localizing the target objects is also vital for video object segmentation tasks. Both RTS[38] and SiamMask[39] demonstrated that pixel-wise masks can provide a more accurate estimation of bounding boxes for VOT tasks. Additionally, their tracking schemes improve the robustness of VOS methods and contribute to target localization in long-term videos.
In 360VOTS, we take advantage of eBFoV to localize the targets on 360°$\degree$ videos and obtain less-distorted search regions. It allows us to exploit arbitrary local visual trackers, which are trained on normal perspective images to achieve enhanced performance on 360°$\degree$ object tracking and segmentation.
@@ III-A Representation for the Target Location
where,
$\small\mathcal{R}_{y}(clon)$ $=\begin{bmatrix}cos(clon)&0&sin(clon)\\ 0&1&0\\ -sin(clon)&0&cos(clon)\end{bmatrix},$ (2)
$\mathcal{R}_{x}(clat)$ $=\begin{bmatrix}1&0&0\\ 0&cos(clat)&-sin(clat)\\ 0&sin(clat)&cos(clat)\end{bmatrix},$ (3)
$\mathcal{R}_{z}(\gamma)$ $=\begin{bmatrix}cos\gamma&-sin\gamma&0\\ sin\gamma&cos\gamma&0\\ 0&0&1\end{bmatrix},$ (4)
@@ IV-A Collection and Annotation
SA The 360°$\degree$ images have stitching artifacts and they affect the target object.
@@
LD The target suffers large distortion due to the equirectangular projection.
@@ IV-B Attribute
The detailed meaning of each attribute is described in Table II. Among them, IV, BC, DEF, MB, CM, ROT, POC, and LD attributes are manually labeled, while the others are computed from the annotation results of targets. The distinct features of the 360°$\degree$ image are well represented in 360VOTS: location variations (FMS, LFoV, and LV), external disturbances (SA and LD) and special imaging (CB and HL).
@@
(a) The category tree of 360VOT.
(b) The category tree of 360VOS.
Figure 6: The tree maps of the category structures for 360VOTS. Newly introduced classes in 360VOS are highlighted in red, while the classes that have been re-categorized from the original 360VOT are highlighted in blue.
@@ IV-C Dataset Statistics of 360VOTS
360VOTS cover four primary categories of humans, animals, rigid objects, and human & carrier cases. As shown in Figure 6, the primary classes are depicted on the branches, with the leaves denoting the respective sub-classes within each primary class. In the majority of sequences, a single target is annotated across all frames, while up to two components of a single target (e.g. bicyclist and bicycle) are annotated separately in several sequences under the human & carrier cases class. Overall, the 120 sequences in the 360VOT dataset are grouped into 32 sub-classes and used for evaluation only. The 360VOS dataset has 62 sub-classes and covers more expansive sequences. In addition, we split the 360VOS dataset into training and test sets with balanced classes of tracking targets, avoiding class bias between the two sets. Eventually, the test set of 360VOS comprises 98,487 annotated frames across 120 sequences and the training set includes 143,634 annotated frames within 170 sequences. As a benchmark suitable for long-term tracking and segmentation evaluation, 360VOS is characterized by an average of 821 frames per sequence in the test set and 845 frames per sequence in the training set, respectively, as reported in Table I.
@@ V-A Evaluation Metrics
$\mathfrak{W}_{u,v}=\|\int_{\varphi-\Delta\varphi/2}^{\varphi+\Delta\varphi/2}\int_{\vartheta-\Delta\vartheta/2}^{\vartheta+\Delta\vartheta/2}\sin\vartheta\;d\vartheta d\varphi\|,$ (13)
where $[\vartheta,\varphi]$ is the longitude and latitude of the spherical coordinate which can be computed from $[u,v]$ via Eq. 6, $\Delta\vartheta=\pi/H$ and $\Delta\varphi=2\pi/W$.
@@ V-B VOT Evaluation
VOT baseline trackers. We evaluated 20 state-of-the-art visual object trackers on 360VOT, as shown in Table III. According to the latest development of visual tracking, the compared methods can be roughly classified into three groups: transformer trackers, Siamese trackers, and other deep learning-based trackers. Specifically, the transformer trackers contain Stark [44], ToMP [45], MixFormer [46], SimTrack [47] and AiATrack [48]. The Siamese trackers include SiamDW [49], SiamMask [39], SiamRPNpp [4], SiamBAN [50], AutoMatch [51], Ocean [52] and SiamX [37]. For other deep trackers, UDT [53], Meta-SDNet [54], MDNet [2], ECO [3], ATOM [55], KYS [56], DiMP [5], PrDiMP [57] were evaluated. We used the official implementation, trained models, and default configurations to ensure a fair comparison among trackers. In addition, we developed a new baseline AiATrack-360 that combines the transformer tracker AiATrack [48] with our 360 tracking framework. We also adapted a different kind of tracker SiamX [37] with our framework, named SiamX-360, to verify the generality of the proposed framework.
@@
(d) (a) Crossing Border
(e) (b) Fast Motion on the Sphere
(f) (c) Latitude Variation
Figure 8: Comparing BBox tracking performances in 360VOT of different trackers, in terms of dual success rate and angle precision rate under the three distinct attributes of 360VOT.
@@
Attribute-based performance on BBox. Furthermore, we evaluated all trackers under 20 attributes to analyze different challenges. In Figure 8, we plot the results on the videos with cross border (CB), fast motion on the sphere (FMS), and latitude variation (LV) attributes. These are the three exclusive and most challenging attributes of 360VOT. Compared to the overall performance, all trackers suffer performance degradation, especially on the CB and FMS attributes. For example, Pdual of SimTrack decreases by 4.2% and 5.3% on CB and FMS respectively. However, the performance of AiATrack-360 still dominates on all three adverse attributes, while SiamX-360 also obtains stable performance gains.
@@ V-C VOS Evaluation
STM[6] 0.364 0.438 0.351 0.426
UNICORN[7] 0.339 0.469 0.328 0.461
TarVis[61] 0.325 0.411 0.320 0.407
AFB-URR[63] 0.389 0.481 0.375 0.472
GMVOS[58] 0.431 0.522 0.416 0.508
CFBI[8] 0.412 0.513 0.397 0.502
CFBI+[64] 0.428 0.532 0.414 0.525
DeAOT[9] 0.447 0.578 0.433 0.569
AOT[62] 0.471 0.597 0.457 0.589
LWL[65] 0.469 0.567 0.457 0.559
TBD[66] 0.474 0.598 0.459 0.590
JOINT[67] 0.487 0.587 0.476 0.581
RTS[38] 0.540 0.645 0.530 0.641
STCN[59] 0.550 0.667 0.535 0.657
XMem[10] 0.560 0.662 0.550 0.659
XMem++[60] 0.579 0.692 0.570 0.687
Xmem 0.596 0.703 0.584 0.699
XMem-360 0.656 0.780 0.639 0.771
XMem-360 0.665 0.786 0.651 0.779
Training details of XMem. We adhered to a similar training scheme of the original XMem[10] to retrain the weights of the XMem network. We first utilized its initialized weights that were pre-trained on static images and BL30K dataset[68], and then we directly performed the main training stage on the combination of our 360VOS training set, YouTubeVOS[19] and DAVIS[18]. To reduce memory cost while learning details on the high-resolution 360°$\degree$ images of 360VOS, we downsampled each frame from 4K resolution to 720p (i.e., 1440$\times$720) for training. During training, all training frames were randomly cropped to $512\times 512$. We employed bootstrapped cross entropy loss and dice loss functions to optimize the model. The optimizer we used is AdamW while the learning rate is set to 1e-5 with a weight decay rate of 0.05. The retraining process lasted around 3.5 days, involving 100K iterations with a batch size of 8 on 4 RTX3090 GPUs.
GT
XMem
XMem
XMem-360
XMem-360
Performance on 360VOS. In each testing sequence, the mask was initialized on the first frame, and all trackers subsequently inferred masks for the following frames. According to the quantitative results reported in Table V, memory-based methods such as STCN[59], XMem[10], and XMem++[60] outperform other existing trackers when dealing with omnidirectional scenes. Particularly, XMem++ [60] achieve the top performance among the 16 existing trackers, with scores of 0.579, 0.692, 0.570, and 0.687 in terms of $\mathcal{J}$, $\mathcal{F}$, $\mathcal{J}_{\text{sphere}}$, and $\mathcal{F}_{\text{sphere}}$, respectively. By leveraging the 360VOS data knowledge, the retrained baseline XMem still show a noticeable improvement of 1.7%, 1.1%, 1.4%, and 1.2% compared to XMem++[60] in terms of $\mathcal{J}$, $\mathcal{F}$, $\mathcal{J}_{\text{sphere}}$, and $\mathcal{F}_{\text{sphere}}$. With our proposed framework, both XMem-360 and XMem-360 demonstrate significant improvements over all other trackers. Notably, XMem-360 and XMem achieve slightly better performance than XMem-360 and XMem, respectively, which present the effectiveness of the 360VOS training set in addressing the intricate appearance variance of tracking targets in 360VOS tasks, as exemplified in Figure 10.

### v2

@@ I Introduction
Furthermore, we present 360VOTS, a challenging benchmark dataset to facilitate research and evaluation on omnidirectional tracking and segmentation, including 360VOT [23] and 360VOS datasets. 360VOT provides dense (rotated) BBox and BFoV annotations of 120 test sequences. Inheriting parts of sequences from the 360VOT dataset, 360VOS incorporates additional 191 sequences and 30 tracking object categories. It contains a total of 290 sequences in 62 categories, while 170 of them are assigned as training set. Importantly, 360VOS provides dense pixel-wise annotations as ground truth that are further utilized to obtain unbiased (rotated) BBoxes and BFoVs for VOT tasks. To accurately evaluate the performance of state-of-the-art (SOTA) approaches on omnidirectional tracking and segmentation, we develop new metrics tailored for 360°$\degree$ images. These metrics encompass dual success, dual (normalized) precision, and angle precision to assess tracking performance, as well as spherical region similarity and spherical contour accuracy to evaluate segmentation performance. Moreover, we exploit the 360VOS training set to retrain the tracker and apply our 360 tracking framework to achieve new SOTA results. With the introduction of new representations, datasets, metrics, and baselines, we believe 360VOTS serves as a vital resource to foster development in omnidirectional visual object tracking and segmentation.
@@
  • We benchmark 24 tracking algorithms and 16 segmentation algorithms on 360VOT and 360VOS respectively, and develop new baselines for future comparisons.
@@ II-A Benchmarks for Visual Object Tracking
ALOV300 [24] is a sparse benchmark introducing 152K frames and 16K annotations. TrackingNet [31] is a large-scale sparse dataset collecting more than 14M frames based on the YT-BB dataset [34]. OxUvA [30] targets long-term tracking by constructing 337 video sequences.
One of the first dense BBox benchmarks is OTB100 [11] which is extended from OTB50 [35] and has 100 sequences. NUS-PRO [25] takes the feature of occlusion-level annotation and provides 365 sequences, while TC128 [26] researches the chromatic information in visual tracking. UAV123 [12], DTB70 [27], and UAVDT [29] offer hundreds of aerial videos containing rigid objects and humans in various scenes. NfS [28] consists of more than 380K frames captured at 240 FPS studying higher frame rate tracking, while LaSOT [14] is a large-scale and category-balanced benchmark of premium quality. GOT-10k [15] provides about 1.5M annotations and 84 classes of objects, aiming at generic object tracking. The annual tracking challenge VOT [13] offered 62 sequences and 20K frames in 2022. Under specific conditions, TOTB [32] focuses on transparent object tracking, while TREK-150 [33] introduces First Person Vision (FPV) videos that interact with the target object.
@@ II-B Benchmarks for Video Object Segmentation
There are several benchmarks for video object segmentation proposed in the last decade as described in Table I. FBMS-59[16] acknowledges the importance of motion information and concentrates on motion segmentation, providing a total of 13,860 frames across 59 sequences. DAVIS[18] is the first dataset designed for video object segmentation (VOS), annotating 376 classes of objects on each frame of 120 short videos. Leveraging the diverse scenarios and photography conditions on YouTube videos, YouTube-VOS[19] introduces a large-scale VOS dataset with more than 540K frames and 3471 training sequences. One of the most recent datasets is LVOS[17] which is the first long-term VOS benchmark dataset with an average length of 560 densely labeled frames. In contrast, our 360VOS provides pixel-wise masks of omnidirectional images for both training and testing.
II-C Other Omnidirectional Vision Benchmarks
In addition to the video-based benchmarks, other related benchmarks such as 360-Indoor [36] and PANDORA [22] are proposed for detection tasks on static 360°$\degree$ images. Meaningful progress has been made by introducing innovative network architecture [37] to address the challenge of large distortions, along with tailored solutions for specific real-world applications [38]. Like omnidirectional data, panoramic data covers a wide FoV, usually up to 360°$\degree$ horizontally. 2D-3D-S [39] and SynPASS [40] are synthetic panoramic benchmarks containing over 70K and 9K RGB images with semantic annotations. Existing methods [41, 40, 42] take advantage of distortion awareness and achieve large advancements in the panoramic semantic segmentation task. Our 360VOS dataset includes numerous high-resolution frames and annotated object-based masks, demonstrating its potential for related tasks through further processing and extension.
II-D Adapting Object Tracking Schemes in Segmentation
To guarantee high tracking speed, the trackers for single object tracking generally crop the image and search for the target in small local regions. The tracking scheme is vital in selecting searching regions and interpreting network predictions over sequences, which enhances tracking performance in the inference phase. For example, DaSiamRPN [43] explored a local-to-global searching strategy for long-term tracking. SiamX [44] proposed an adaptive inference scheme to prevent tracking loss and realized fast target re-localization. Similarly, localizing the target objects is also vital for VOS tasks. Both RTS[45] and SiamMask[46] demonstrated that pixel-wise masks can provide a more accurate estimation of bounding boxes for VOT tasks. Additionally, their tracking schemes improve the robustness of VOS methods and contribute to target localization in long-term videos.
In 360VOTS, we utilize eBFoV to localize the targets on 360°$\degree$ videos and obtain less-distorted search regions. It allows us to adopt arbitrary local visual trackers trained by perspective images on 360°$\degree$ object tracking and segmentation.
@@ III-A Representation for the Target Location
where $\mathcal{R}$ denotes the 3D rotation along the $y$, $x$, $z$ axes,
@@
The 360 tracking framework comprises three essential hyper-parameters: SR Ratio, SR Min, and Max Loss. When utilizing the previously predicted target mask, we extend the search region in the subsequent frame by a scaling factor known as SR Ratio. Moreover, SR Min regulates the minimum BFoV of the search region, preventing the system from becoming fixated on excessively small search areas, particularly when the tracking target is tiny. Once the target is lost, the framework maintains the search region from the last successful frame for a fixed number of consecutive frames (i.e., Max Loss), allowing the tracker to relocate the target within the same area. If the target remains lost beyond this threshold, the search region is progressively expanded to account for the potential displacement. This expansion continues until the search region either covers the entire 360°$\degree$ or a predefined limit on consecutive frames. If the target remains undetected beyond this stage, the framework initiates a full-frame search, ensuring recovery in extreme cases. Once the target is re-tracked, the search region is dynamically updated based on the updated target BFoV, allowing the framework to seamlessly resume normal tracking operations.
@@ IV-A Collection and Annotation
SA The images have stitching artifacts affecting the target object.
@@
LD The target suffers large distortion due to equirectangular projection.
@@ IV-B Attribute
The detailed meaning of each attribute is described in Table II. Among them, IV, BC, DEF, MB, CM, ROT, POC, SA, and LD attributes are manually labeled, while the others are computed from the annotation results of targets. The distinct features of the 360°$\degree$ image are well represented in 360VOTS: location variations (FMS, LFoV, and LV), external disturbances (SA and LD) and special imaging (CB and HL).
@@ IV-C Dataset Statistics of 360VOTS
360VOTS cover four primary categories of humans, animals, rigid objects, and human & carrier cases. The sub-classes within each primary class are depicted in the supplementary. In the majority of sequences, a single target is annotated across all frames, while up to two components of a single target (e.g. bicyclist and bicycle) are annotated separately in several sequences under the human & carrier cases class. Overall, the 120 sequences in the 360VOT dataset are grouped into 32 sub-classes and used for evaluation only. The 360VOS dataset has 62 sub-classes and covers more expansive sequences. In addition, we split the 360VOS dataset into training and test sets with balanced classes of tracking targets, avoiding class bias between the two sets. Eventually, the test set of 360VOS comprises 98,487 annotated frames across 120 sequences and the training set includes 143,634 annotated frames within 170 sequences. As a benchmark suitable for long-term tracking and segmentation evaluation, 360VOS is characterized by an average of 821 frames per sequence in the test set and 845 frames per sequence in the training set, respectively, as reported in Table I.
@@ V-A Evaluation Metrics
$\mathfrak{W}_{u,v}=\|\int_{\theta-\Delta\theta/2}^{\theta+\Delta\theta/2}\int_{\phi-\Delta\phi/2}^{\phi+\Delta\phi/2}\sin\phi\;d\phi d\theta\|,$ (10)
where $[\theta,\phi]$ is the longitude and latitude of the spherical coordinate which can be computed from $[u,v]$ via Eq. 3, $\Delta\theta=2\pi/W$ and $\Delta\phi=\pi/H$.
@@ V-B VOT Evaluation
VOT baseline trackers. We evaluated 24 state-of-the-art visual object trackers on 360VOT, as shown in Table III. The compared methods can be roughly classified into three groups: transformer trackers, Siamese trackers, and other deep learning-based trackers. Specifically, the transformer trackers contain ARTrack [51], HIPTrack [52], OSTrack [53], LoRAT [54], Stark [55], ToMP [56], MixFormer [57], SimTrack [58] and AiATrack [59]. The Siamese trackers include SiamDW [60], SiamMask [46], SiamRPNpp [4], SiamBAN [61], AutoMatch [62], Ocean [63] and SiamX [44]. For other deep trackers, UDT [64], Meta-SDNet [65], MDNet [2], ECO [3], ATOM [66], KYS [67], DiMP [5], PrDiMP [68] were evaluated. We used the official implementation, trained models, and default configurations to ensure a fair comparison among trackers. In addition, we developed a new baseline AiATrack-360 that combines the transformer tracker AiATrack [59] with our 360 tracking framework. We also adapted a different kind of tracker SiamX [44] with our framework, named SiamX-360, to verify the generality of the proposed framework.
BBox
rBBox
BFoV
rBFoV
@@ V-C VOS Evaluation
STM[6] 0.364 0.438 0.366 0.439
UNICORN[7] 0.339 0.469 0.340 0.470
TarVis[72] 0.324 0.411 0.325 0.414
AFB-URR[74] 0.389 0.481 0.392 0.483
GMVOS[69] 0.431 0.522 0.435 0.526
CFBI[8] 0.412 0.513 0.414 0.516
CFBI+[75] 0.428 0.532 0.430 0.534
DeAOT[9] 0.447 0.578 0.448 0.579
AOT[73] 0.471 0.597 0.473 0.598
LWL[76] 0.469 0.567 0.471 0.569
TBD[77] 0.474 0.598 0.477 0.600
JOINT[78] 0.487 0.587 0.488 0.588
RTS[45] 0.540 0.645 0.542 0.647
STCN[70] 0.550 0.667 0.552 0.669
XMem[10] 0.560 0.662 0.562 0.663
XMem++[71] 0.579 0.692 0.581 0.693
Xmem 0.596 0.703 0.597 0.704
XMem-360 0.656 0.780 0.658 0.782
XMem-360 0.676 0.799 0.677 0.801
@@
GT
XMem
XMem
XMem-360
XMem-360
Performance on 360VOS. In each testing sequence, the mask was initialized on the first frame, and all trackers subsequently inferred masks for the following frames. According to the quantitative results reported in Table V, memory-based methods such as STCN[70], XMem[10], and XMem++[71] outperform other existing trackers when dealing with omnidirectional scenes. Particularly, XMem++ [71] achieve the top performance among the 16 existing trackers, with scores of 0.579, 0.692, 0.570, and 0.687 in terms of $\mathcal{J}$, $\mathcal{F}$, $\mathcal{J}_{\text{sphere}}$, and $\mathcal{F}_{\text{sphere}}$, respectively. With our proposed framework, both XMem-360 and XMem-360 demonstrate significant improvements over all other trackers. The visualized results are presented in Figure 8.
(a) Search Region Ratio
(b) Min Search Region
(c) Max Frames Loss
Figure 9: Ablation studies of the 360 tracking framework. The baseline, represented by the black dot in each plot, corresponds to XMem-360 with specific settings. The performance is measured by average spherical metrics $(\mathcal{J}\&\mathcal{F})_{sphere}$. Other observations in each setting group are marked in different colored dots. The tendency lines are computed using the least squares method based on the baseline and all observations.
@@
Robustness of 360 tracking framework. We have conducted ablation experiments to analyze the impact of different settings on our proposed 360 tracking framework in Section III-B, including 3 key parameters: SR Ratio, SR Min, and Max Loss. As demonstrated in Figure 9, the performance variations are within a narrow range (e.g., ±$\pm 5\%$) apart from extreme cases, showing that our framework is stable and robust across a wide range of parameter values. The baseline setting provides the best trade-off between accuracy and robustness. The extreme settings (e.g., SR Min = 60°), result in a larger drop in performance (e.g., $-6.8\%$) due to overly restrictive search regions. Max Loss indicates how many frames the search region remains unchanged after the tracker loses the target. Disabling this mechanism entirely (e.g., Max Loss = 0) results in more significant performance degradation (e.g., $-5.1\%$), highlighting the importance of the search region updating strategy in our framework.
V-D Efficacy of 360VOTS Training Sets
360VOTS fills in the absence of training data for omnidirectional tracking, as mentioned in Section IV-C. Firstly, we retrained XMem [10] on a combined dataset that consists of 360VOTS training sets, YouTubeVOS[19], and DAVIS[18]. According to Table V, the retrained baseline XMem show a noticeable improvement of 1.7%, 1.1%, 1.4%, and 1.2% compared to XMem++[71] in terms of $\mathcal{J}$, $\mathcal{F}$, $\mathcal{J}_{\text{sphere}}$, and $\mathcal{F}_{\text{sphere}}$. After integrating the retrained model with our 360 tracking framework, XMem-360 achieves the best performance. The new training data enhances models in addressing the intricate appearance variance of tracking targets in 360VOS, illustrated in Figure 8. Besides retraining a VOS tracker, we converted the segmentation masks to obtain VOT ground truth and fine-tuned the pretrained model of LoRAT[54] on 360VOTS training sets only for 10 epochs. The finetuned VOT tracker LoRAT achieves 0.495, 0.504, 0.511, and 0.526 in terms of Sdual, Pdual, $\overline{\mbox{P}}_{dual}$, and Pangle in 360VOT, which outperforms LoRAT[54] 7.38%, 7.69%, 4.93%, and 4.57% accordingly. Even though only conducted rough finetuning, LoRAT shows competitive performance compared to AiATrack-360. The consistent advancements of XMem and LoRAT verify the efficacy of training data in both 360VOS and 360VOT, matching the objectives on the newly proposed datasets.